Four point charges are equally separated by 1.00 m in air. If q1 = 3.0 nC, q2 = -4.0 nC, q3 = 2.0 nC, and q4 = -5.0 nC, what is the electric field at a point halfway between q2 and q3? show the equation, solution and complete answer.


1
Expert's answer
2022-01-31T10:13:53-0500

The electric field at a point halfway between q2 and q3 is


"E=\dfrac{1}{4\pi\epsilon_0}\bigg(\dfrac{q_1}{(r+r/2)^2}+\dfrac{q_2}{(r/2)^2}-\dfrac{q_3}{(r/2)^2}-\dfrac{q_4}{(r+r/2)^2}\bigg),\\\space\\
E=-183.7\text{ V/m}
."


The negative sign means that the field points to the left.